5. The school that Chris goes to is selling tickets to a fall musical. Prices are $6 for adult tickets and $5 for child tickets. The school took in $500 total. They sold 88 tickets. How many adult tickets and how many child tickets were sold?


Answer by macston(5194)

A= adult; C=child
A+C=88
6A+5C=$500
Solve first equation for A
A+C=88 subtract C from each side
A+C-C=88-C
A=88-C
Substitute for A in second equation
6(88-C)+5C=$500
528-6C+5C=500
528-C=500 add C to Each side
528-C+c=500+C Subtract 500 from each side
528-500=100+C-500
28=C ANSWER 1:28 Child tickets were sold
A=88-C=88=28=60 ANSWER 2:60 Adult tickets were sold
CHECK:
$6A+&5C=$500
$6(60)+$5(28)=$500
$360+$140=$500
$500=$500